# Read Your Own Mind: Reasoning Helps Surface Self-Confidence Signals in LLMs

### v1

@@ Appendix A Data Creation and Composition
Each example received two human labels - Knowledge Domain, and Skill Required. A large language model (OpenAI o3) proposed initial tags for 100 random questions. The first author then reviewed every instance, correcting tags where needed, and used these tags to manually label all 270 samples. You can find the specific tags and number of datapoints in Figure 5. In the main paper, we break out results for the full dataset and for the two most common Skill Required tags only; Knowledge Domain splits are omitted because several categories are too small. Per-dataset results can be found in Appendix C. Five representative questions and their tags are shown in Table 2.
@@
In what year did Augustus De Morgan publish the article ”Trochoidal Curve” in the Penny Cyclopaedia? SimpleQA Fact Retrieval History and Past Events
@@
Appendix B Prompts and Inference
@@
[…]
@@
    Multiple-choice reformulation. We label the distinct candidates with letters A, B, C…, we also add an option ”None  Unknown”.
@@
“You are a helpful assistant. We asked a person to answer an open-ended question. The person wrote a reasoning trace and then gave a final answer. I want to know what all the possible options were that the person considered before giving the final answer. This is for research on the answers people consider. I want to use these possible answers to create a multiple-choice question. Please give me all the possible answers the person considered, without duplicates, keeping the answers distinct and suitable for a multiple-choice question. Respond as a JSON list of strings, like this: FINAL LIST: [”answer1”, ”answer2”, …], which I can later use in a quiz. Add the correct answer, the final answer, and Other / Unknown to the list, making sure there are no duplicates. Reason about possible answers and how to remove duplicates, before giving the final list. Include only answers mentioned in the reasoning trace plus the correct answer. Do not add answers that are missing from the trace or the correct / unknown items.”
@@
Appendix C Results Across Datasets

### v2

@@ Appendix B Data Creation and Composition
Each example received two human labels - Knowledge Domain, and Skill Required. A large language model (OpenAI o3) proposed initial tags for 100 random questions. The first author then reviewed every instance, correcting tags where needed, and used these tags to manually label all 270 samples. You can find the specific tags and number of datapoints in Figure 5. In the main paper, we break out results for the full dataset and for the two most common Skill Required tags only; Knowledge Domain splits are omitted because several categories are too small. Per-dataset results can be found in Appendix D. Five representative questions and their tags are shown in Table 1.
@@
In what year did Augustus De Morgan publish the article "Trochoidal Curve" in the Penny Cyclopaedia? SimpleQA Fact Retrieval History and Past Events
@@
Appendix C Prompts and Inference
@@
    Candidate extraction. For each question, we feed the entire $3.4\text{k}$-token reasoning chain to gpt-4o-mini, prompting it to list all candidate answers mentioned in the trace.
@@
    Multiple-choice reformulation. We label the distinct candidates with letters A, B, C…, we also add an option "None Unknown".
@@
“You are a helpful assistant. We asked a person to answer an open-ended question. The person wrote a reasoning trace and then gave a final answer. I want to know what all the possible options were that the person considered before giving the final answer. This is for research on the answers people consider. I want to use these possible answers to create a multiple-choice question. Please give me all the possible answers the person considered, without duplicates, keeping the answers distinct and suitable for a multiple-choice question. Respond as a JSON list of strings, like this: FINAL LIST: ["answer1", "answer2", …], which I can later use in a quiz. Add the correct answer, the final answer, and Other / Unknown to the list, making sure there are no duplicates. Reason about possible answers and how to remove duplicates, before giving the final list. Include only answers mentioned in the reasoning trace plus the correct answer. Do not add answers that are missing from the trace or the correct / unknown items.”
@@
Appendix D Results Across Datasets